OpenStudy (anonymous):

Find an equation of the parabola that has the indicated vertex and whose graph passes through the given point Vertex: ( 2, -1) ; point: ( 4, -3)

OpenStudy (mathmate):

Do you know if the parabola has a vertical or horizontal axis?

OpenStudy (anonymous):

no I don't

OpenStudy (mathmate):

Do you expect it to look like (y-k)=a(x-h)^2, or (x-h)=a(y-k)^2? 

OpenStudy (anonymous):

I am not sure what it is suppose to look like....I just need to find an equation....I don't know how to solve the problem

OpenStudy (mathmate):

There are two solutions. The equation could be of the form (y-k)=a(x-h)^2, or (x-h)=a(y-k)^2 where (h,k) is the location of the vertex (2,-1), and a is a parameter to be found.

OpenStudy (anonymous):

Thank you but how do you work out the problem

OpenStudy (mathmate):

Substitute (h,k) into the equations, and the given point into the x- and y-values to solve for a. You can choose one of the two forms, or do both.

OpenStudy (anonymous):

Where do I put the point values of (4, -3) in the equation?

OpenStudy (mathmate):

For example, if a parabola has a vertex (3,5) and passes through (5,9), then h=3, k=5, and put x=4 and y=9 to find a. For vertical axis, we use (y-k)=a(x-h)^2 to get (9-5)=a(5-3)^2 solving for a gives a=1, so the equation is (y-5)=(x-3)^2

OpenStudy (mathmate):

* x=5

OpenStudy (anonymous):

how did u get a=1?

OpenStudy (mathmate):

(9-5)=a(5-3)^2 4=a(2^2) 4=4a a=1